OpenStudy (moonlitfate):

Use implicit differentiation to determine dy/dx if x*e^y+1=x*y

OpenStudy (e.mccormick):

OK. Do you know how implicit works?

OpenStudy (moonlitfate):

A bit, yes. I know it's basically chain ruling.

OpenStudy (e.mccormick):

Yah. The whole goal is that wherever you have a y it becomes a dy/dx, then you solve for that. Now, this isa little odd in the question part so... Are you trying to determine if the left of the = is the same as the right?

OpenStudy (e.mccormick):

OR, it set as x*e^y+1=x*y and you want to do the whole thing. I have seen these things go both ways in questions.

OpenStudy (e.mccormick):

From the way you wrote it, I am betting the second... which will be fun because it means you have to differentiate both sides of the equation.

OpenStudy (moonlitfate):

Yeah, both sides, lol.

OpenStudy (e.mccormick):

OK, so know the setup? You are looking at product rules on both sides with a chained on dy every time you differentiate a y.

OpenStudy (moonlitfate):

So it would dy/dx(x*e^y)+dy/dx[1]=dy/dx[x*y]

OpenStudy (moonlitfate):

e^y*x+0=xy'

OpenStudy (moonlitfate):

Yeah, I know I used two different notations, lol. y' is easier to type.

OpenStudy (e.mccormick):

np

OpenStudy (e.mccormick):

The right side would be xy, so product rule on differentiation.

OpenStudy (e.mccormick):

Actually, product rule on both sides.

OpenStudy (e.mccormick):

\[\frac{d}{dx}[xe^y+1]=\frac{d}{dx}[xy]\Rightarrow \frac{d}{dx}[xe^y]+ \frac{d}{dx}[1]=\frac{d}{dx}[xy]\Rightarrow \frac{yd}{dx}[xe^y]+1e^y + 0=\frac{yd}{dx}x+1y\]

OpenStudy (e.mccormick):

Oops. Meant dy/dx, not yd. LOL!

OpenStudy (moonlitfate):

Lol, np. x3

OpenStudy (e.mccormick):

So this, before it is solved for \(\frac{dy}{dx}\) \[\frac{dy}{dx}[xe^y]+e^y =\frac{dy}{dx}x+1y\]

OpenStudy (e.mccormick):

Now, there are two solutions from that point, but they differ very little. It depends on which side you pull the \(\frac{dy}{dx}\) to.

OpenStudy (e.mccormick):

Do you get what I am taling about there?

OpenStudy (moonlitfate):

Yes.

OpenStudy (e.mccormick):

OK, so where would you go with it from that point?

OpenStudy (moonlitfate):

all the dy/dx will be on th eleft side.

OpenStudy (e.mccormick):

OK. So show me your version. I'm doing both. Hehe.

OpenStudy (moonlitfate):

Lol.

OpenStudy (moonlitfate):

dy/dx = -e^y+y/(xe^y-x) ? o:

OpenStudy (e.mccormick):

Any luck over there?

OpenStudy (moonlitfate):

I just typed out an answer.

OpenStudy (e.mccormick):

Hehe. About the same time. LOL. \[e^y-1y =\frac{dy}{dx}x-\frac{dy}{dx}xe^y\\ e^y-1y =\frac{dy}{dx}(x-xe^y)\\ \frac{e^y-1y}{x-xe^y} =\frac{dy}{dx}\] \[\frac{dy}{dx}xe^y-\frac{dy}{dx}x =1y-e^y\\ \frac{dy}{dx}(xe^y-x) =1y-e^y\\ \frac{dy}{dx} =\frac{1y-e^y}{xe^y-x}\]

OpenStudy (moonlitfate):

Haha, nice. :D

OpenStudy (e.mccormick):

Hehe. Thought you might like to see the way I did it. I mean, it is just algebra, but you really have to watch those sign changes! 90% of calculus is algebra, trig, and everything down to 1+1=2.

OpenStudy (e.mccormick):

In derivatives, as you know, you do things like \(nx^{n-1}\). When you get to anti-derivatives (integrals) you will do this in reverse. If you think about it, \(n\pm 1\) can be \(2-1\) and \(1+1\), AKA: basic arithmetic. So back in first grade, you were learning skills required to do calculus and they never bothered to tell you!